OpenStudy (anonymous):

help me solve this please 3x-4 / 4-3x + 6-x / x-6 = 0 I know you find the Domain first which is 6, 4/3 The answer is no solution, but how do you do it?

OpenStudy (anonymous):

\[\frac{3x-4}{4-3x}+\frac{6-x}{x-6} = 0\] \[\frac{(3x-4)(x-6) + (6-x)(4-3x)}{(4-3x)(x-6)} = 0\] \[(3x-4)(x-6)+(6-x)(4-3x)= 0\]

OpenStudy (anonymous):

\[3x ^{2}-22x+24+24-22x+3x ^{2}\]

OpenStudy (anonymous):

\[6x ^{4}-44x ^{2}+48=0? I feel like I totally messed up?\]